OpenStudy (anonymous):

What is the solution set of the equation 2y^2 - 3y = 2 The solution set is {-1/2, [1]}

OpenStudy (freckles):

this is a quadratic equation you can subtract 2 on both sides and put it in the form: ay^2+by+c=0

OpenStudy (anonymous):

I'm still very confused. Would you be able to tell me exactly how to do it?

OpenStudy (freckles):

Are you asking me how to subtract 2 on both sides?

OpenStudy (freckles):

2y^2-3y-2=2-2 2y^2-3y-2=0

OpenStudy (freckles):

now you can try to factor or use quadratic formula